Some cell membranes in the human body have a layer of negative charge on the inside surface and a layer of positive charge of equal magnitude on the outside surface. Suppose that the charge density on either surface is ± 0.50×10−3 C/m2, the cell membrane is 5.00 nm thick, and the cell membrane is made up of tissue with a dielectric constant of 5.40.

A) Find the magnitude of E in the membrane between the two layers of charge.

E=? V/m

B) Find the potential difference between the inside and the outside of the cell. Which is at the higher potential?

ΔV=?V

C) A typical cell in the human body has a volume of 10−16m3 . Estimate the total electric-field energy stored in the wall of a cell of this size. (Hint: The cell membrane is very thin. Assume that the cell is spherical, and calculate the area of the cell membrane.)

U=?J

I am really having trouble with this problem; nothing i try is correct. I appreciate the help.
Answer #1


Electric field between two surfaces can be calculated by using E = keo Here, is the electric charge density, kis the dielectr

Substitute 0.5x 103 C/m2fora,5.4for k and 8.85 x1012 F/m 0.5x10-3 C/m2 E= (5.4) (8.85x10-12 F/m) E 10.46 x 106 V/m The value

Volume of cell V = 10-16 mm3 4 V Finding the value ofradius ofthe cell 3 10-16 mm 4T R 2.87x105 Volume of the shell mm shell